# Zero-shot OCR Accuracy of Low-Resourced Languages: A Comparative Analysis on Sinhala and Tamil

### v1

@@ 3.3 Integration of OCR Systems
The process of integrating the Tesseract engine was quite simple through the use of the pytesseract library. This library enabled the automation of character recognition for every record within each dataset, streamlining the overall data processing workflow.
As previously mentioned, the Subasa OCR engine is exclusively accessible via a web application. This limitation meant that we had to manually input images one by one to perform OCR, a process that proves impractical given our substantial Sinhala dataset of 6,969 records. To streamline this tedious task and enhance efficiency, we resolved to automate it using Selenium after meticulously identifying the element locations by inspecting the web source code. This strategic shift not only alleviates manual effort but also drastically reduces processing time.
The seamless integration of Surya and EasyOCR was straightforward, largely due to their excellent documentation. Both engines are accessible on GitHub and can be conveniently installed as libraries directly through Python, making the setup process efficient and user-friendly.
@@ 3.4 Evaluation Mechanism
The evaluation was done using the generated text against the reference text of the datasets. For the comparison, we employed five different measurements. They are listed as below.
@@
Word Error Rate (WER): Similar to the CER, the WER is calculated by comparing the text produced by the OCR system against a ground truth or reference text. The WER is determined by the number of word-level errors made by the OCR engine. The formula for calculating the word error rate is also $(S+I+D)/N$, but considered at the word level is instead of the character level Nazeem et al. (2024).
@@
Average Normalised Levenshtein Similarity (ANLS): This metric takes into account both reasoning errors and shortcomings of OCR. To evaluate answers, it calculates a similarity score between the model’s response and the ground truth using Levenshtein distance. A key feature of this scoring system is its application of a threshold ($\tau$=0.5) on the normalised Levenshtein distance (NL): if NL is less than or equal to 0.5, the similarity score is calculated as 1-NL; otherwise, the score is 0. This approach allows ANLS to provide intermediate scores (ranging from 0.5 to 1) for responses that are logically correct but may contain minor recognition errors, contrasting with standard accuracy metrics, which would score them as zero Biten et al. (2019).
@@
Metric for Evaluation of Translation with Explicit ORdering (METEOR): METEOR is an metric designed for evaluating the quality of machine translation. It operates by identifying generalised unigram matches between machine-generated outputs and human reference translations, allowing for matches based on surface forms, stemmed forms, and synonyms. Designed specifically to address limitations of the BLEU metric, such as its inability to directly account for recall and its indirect measurement of word order, METEOR computes a score by integrating unigram precision, unigram recall (with greater emphasis on recall), and a fragmentation penalty that evaluates the word order of matched words. This method has demonstrated improved correlation with human judgments when compared to other evaluation techniques. Given its enhanced ability to assess generated text against references and its rapid adoption in OCR studies, we chose to utilise this metric Banerjee and Lavie (2005).
3.5 Discussion of Results
@@
In this section, we present the analysis of the results drawn from our comparative evaluation of OCR systems. The findings are showcased distinctly for both Sinhala and Tamil in Table 1. With the exception of Subasa OCR and EasyOCR, all four other systems were able to process both languages selected for evaluation in this study. It is important to note that Subasa OCR is a monolingual system specifically fine-tuned for Sinhala. While EasyOCR has multilingual capabilities, it notably lacks support for the Sinhala language.
When evaluating the overall results for both languages, the Cloud Vision API and Document AI have achieved very similar results. Notably, Document AI outperforms the other engines for Tamil across all metrics.
Overall, the best WER results for Tamil are notably higher than those for Sinhala. This observation highlights the disparity between character and word identification. Despite achieving a very low CER by Document AI, the comparatively higher WER indicates that while the system effectively identifies characters, it struggles with word formation and spacing of the Tamil language. In contrast, the METEOR and ANLS scores for both languages are relatively high, suggesting a strong alignment in terms of content, word order, and semantic meaning. However, the BLEU scores for Tamil are markedly lower than those of other metrics, likely due to the elevated WER, which results in fewer successful n-gram overlaps.
Surya’s performance in the Sinhala language has been nothing short of extraordinary, emerging as the standout among the others. The metrics clearly illustrate this success, as highlighted in Table 1. When we compare the best WER for Tamil at 11.98% with that of Sinhala at an impressive 2.61% as depicted in Figure 4, the superiority of the Surya engine’s accuracy for Sinhala becomes strikingly apparent. Furthermore, the METEOR and ANLS scores of 0.9723 and 0.9920, respectively, underscore its near-perfect word-level performance. This data powerfully demonstrates Surya’s exceptional capabilities in processing the Sinhala language.
(a) Sinhala
(b) Tamil
Figure 4: WER results for Sinhala and Tamil
The comparison between Subasa OCR and Tesseract is particularly compelling, as Subasa OCR represents a fine-tuned adaptation of the Tesseract 4.0 engine specifically for Sinhala. The authors of Subasa assert that their modifications yield significantly superior results compared to the standard Tesseract 4.0 Anuradha et al. (2021). The metric evaluations reveal that Tesseract 5.5.0 outperforms Subasa OCR across all metrics. This indicates that the latest version of Tesseract by Google has made substantial enhancements for the Sinhala language, even in its vanilla form. However, Tesseract’s performance in Tamil is not competitive compared to other systems.
@@ 4 Conclusion
The results indicated that Document AI performed best for Tamil, while Surya excelled in Sinhala. The overall performance of both Sinhala and Tamil by the Cloud Vision API and Document AI overall is reasonable, highlighting the capabilities of commercial engines in the field of OCR, as anticipated. A standout performer was Surya for Sinhala, which outperformed all other OCR systems in each metric. Furthermore, the significant disparity between the best CER and WER results for Tamil indicates that while the system excels at character recognition, it falls short in accurately identifying words through proper character formation and whitespace detection. Additionally, it was remarkable to find that Zero-Shot Tesseract 5.5.0 outperformed a Tesseract 4.0 system fine-tuned on Sinhala (Subasa OCR).
@@
The analysis conducted focused on comparing the performance of selected OCR engines using Tamil and Sinhala printed scripts. However, it is important to note a limitation regarding the datasets used; both are synthetically created, featuring black text on a white background. This design results in clean and clear images, which do not accurately reflect real-world conditions when capturing printed text. OCR technology faces significant challenges related to the quality of input images, especially when dealing with historical documents or low-resource language data. Factors such as poor print quality, low resolution, shading, blurriness, shine-through effects, stains, and skewness can severely impact OCR accuracy Hegghammer (2022). Images that exhibit distortion, textured backgrounds, cluttered environments, disconnected line segments, isolated dots, breaks in lines, rotation, motion blur, and out-of-focus blur complicate character segmentation and recognition, often leading to higher error rates Karunarathne et al. (2017). Additionally, low image resolutions can hamper overall OCR speed, as uncertainty in character representation results in a greater number of recognition variants Anuradha et al. (2021). Consequently, when evaluated using camera-captured images, the accuracy levels may notably differ from the results presented in this study.

### v2

@@ 3.3 Integration of OCR Systems
The process of integrating the Tesseract engine was quite simple through the use of the pytesseract [^29] https://pypi.org/project/pytesseract/ wrapper. This enabled the automation of character recognition for every record within each dataset, streamlining the overall data processing workflow.
As previously mentioned, the Subasa OCR engine is exclusively accessible via a web application. This limitation meant that we had to manually input images one by one to perform OCR, a process that proves impractical given our substantial Sinhala dataset of 6,969 records. To streamline this tedious task and enhance efficiency, we resolved to automate it using Selenium [^30] https://www.selenium.dev/ after meticulously identifying the element locations by inspecting the web source code.
The seamless integration of Surya and EasyOCR was straightforward, largely due to their comprehensive documentation. Both engines are accessible on GitHub and can be conveniently installed as libraries directly through Python, making the setup process efficient and user-friendly.
@@ 3.4 Evaluation Mechanism
The evaluation was performed using the generated text against the reference text of the datasets. For the comparison, we employed five different measurements;
@@
Word Error Rate (WER): Similar to the CER, the WER is calculated by comparing the text produced by the OCR system against a ground truth or reference text. The WER is determined by the number of word-level errors made by the OCR engine. The formula for calculating the word error rate is also $(S+I+D)/N$, but considered at the word level instead of the character level Nazeem et al. (2024).
@@
Average Normalised Levenshtein Similarity (ANLS): This metric takes into account both reasoning errors and shortcomings of OCR. To evaluate answers, it calculates a similarity score between the response of the model and the ground truth using Levenshtein distance. A key feature of this scoring system is its application of a threshold ($\tau$=0.5) on the normalised Levenshtein distance (NL): if NL is less than or equal to 0.5, the similarity score is calculated as 1-NL; otherwise, the score is 0. This approach allows ANLS to provide intermediate scores (ranging from 0.5 to 1) for responses that are logically correct but may contain minor recognition errors, contrasting with standard accuracy metrics, which would score them as zero Biten et al. (2019).
@@
Metric for Evaluation of Translation with Explicit ORdering (METEOR): Similar to BLEU, METEOR is also a metric designed for evaluating the quality of machine translation. It operates by identifying generalised unigram matches between machine-generated outputs and human reference translations, allowing for matches based on surface forms, stemmed forms, and synonyms. Designed specifically to address limitations of the BLEU metric, such as its inability to directly account for recall and its indirect measurement of word order, METEOR computes a score by integrating unigram precision, unigram recall (with greater emphasis on recall), and a fragmentation penalty that evaluates the word order of matched words. This method has demonstrated improved correlation with human judgments when compared to other evaluation techniques. Given its enhanced ability to assess generated text against references and its rapid adoption in OCR studies, we chose to utilise this metric Banerjee and Lavie (2005).
4 Discussion of Results
@@
[…]
Figure 4: Results for Sinhala and Tamil
In this section, we present the analysis of the results drawn from our comparative evaluation of OCR systems. The findings are showcased distinctly for both Sinhala and Tamil in Table 1. With the exception of Subasa OCR and EasyOCR, all four other systems were able to process both languages selected for evaluation in this study. It is important to note that Subasa OCR is a monolingual system specifically fine-tuned for Sinhala. While EasyOCR has multilingual capabilities, it notably lacks support for the Sinhala language. When evaluating the overall results for both languages, the Cloud Vision API and Document AI have achieved very similar results. However, Document AI outperforms the other engines for Tamil across all metrics.
Overall, the best WER results for Tamil are notably higher than those for Sinhala. This observation highlights the disparity between character and word identification. Document AI, despite achieving a very low CER, achieves a comparatively higher WER, indicating that while the system effectively identifies characters, it struggles with word formation and spacing of the Tamil language. This issue is common across all engines and applies to both languages, as systems tend to perform better at recognising characters but struggle with forming words and managing spacing. In contrast, the METEOR and ANLS scores for both languages are relatively high, suggesting a strong alignment in terms of content, word order, and semantic meaning. However, the BLEU scores for Tamil are markedly lower than those of other metrics, likely due to the elevated WER, which results in fewer successful n-gram overlaps.
Performance of Surya on the Sinhala language has been nothing short of extraordinary, emerging as the standout among the others. The metrics clearly illustrate this success, as highlighted in Table 1. When we compare the best WER for Tamil at 11.98% with that of Sinhala at an impressive 2.61% as depicted in Figure 4, the superiority of the accuracy of the Surya engine for Sinhala becomes strikingly apparent. Furthermore, the METEOR and ANLS scores of 0.9723 and 0.9920, respectively, underscore its near-perfect word-level performance.
The comparison between Subasa OCR and Tesseract is particularly compelling, as Subasa OCR represents a fine-tuned adaptation of the Tesseract 4.0 engine specifically for Sinhala. The authors of Subasa assert that their modifications yield significantly superior results compared to the standard Tesseract 4.0 Anuradha et al. (2021). The metric evaluations reveal that Tesseract 5.5.0 now outperforms Subasa OCR across all metrics. This indicates that the latest version of Tesseract by Google has made substantial enhancements for the Sinhala language, even in its vanilla form. However, Tesseract’s performance in Tamil is not competitive compared to other systems.
@@
Furthermore, we performed a character-level error analysis for the best models of each language. This analysis involved counting the number of errors per character by comparing the generated text to a reference. In Sinhala, we identified erroneous characters based on a threshold of more than 5,000 errors. Diacritics such as ‘ [Uncaptioned image] ’, ‘ [Uncaptioned image] ’, ‘ [Uncaptioned image] ’, and ‘ [Uncaptioned image] ’ were particularly difficult to identify. In addition, letters such as ‘ [Uncaptioned image] ’, ‘ [Uncaptioned image] ’, ‘ [Uncaptioned image] ’, ‘ [Uncaptioned image] ’, and ‘ [Uncaptioned image] ’ posed challenges for Surya in terms of accurate detection.
For Tamil, we set a threshold of more than 1,600 errors to identify erroneous characters. The diacritic ‘ [Uncaptioned image] ’ emerged as the most error-prone character, while letters such as ‘ [Uncaptioned image] ’, ‘ [Uncaptioned image] ’, ‘ [Uncaptioned image] ’, and ‘ [Uncaptioned image] ’ were also among the most problematic for Document AI.
The number of character errors is generally higher for Sinhala compared to Tamil. However, the best CER scores for both languages do not highlight this difference. This discrepancy arises because, despite having a greater number of character-level errors, the overall usage frequency per character in Sinhala is also comparably higher, which reduces the average error rate. This phenomenon is potentially due to the differences in the sizes of their alphabets [^31] The moniker Alphabet is used in the general meaning here. Both these scripts are in fact Abugidas rather than Alphabets.: Sinhala has 60 characters, while Tamil contains 247.
@@ 5 Conclusion
The results indicated that Document AI performed best for Tamil, while Surya excelled in Sinhala. Both the Cloud Vision API and Document AI showed reasonable performance in OCR for these languages, highlighting the capabilities of commercial engines, as anticipated. A standout performer was Surya for Sinhala, which outperformed all other OCR systems in each metric. Furthermore, the significant disparity between the best CER and WER results for Tamil compared to Sinhala indicates that while Document AI excels at character recognition, it falls short in accurately identifying words through proper character formation and white-space detection. Additionally, zero-shot Tesseract 5.5.0 outperformed a fine-tuned Tesseract 4.0 system on Sinhala (Subasa OCR). Moreover, The differences in scores between the commercial OCR systems are largely a black box, likely arising from nuances in their architectures and training data.
Out of approximately 2.2 million Tamil text records, we selected only 7,000 records to ensure a fair comparison with the Sinhala dataset, which has only 6,969 records. In future studies, it would be possible to expand the two synthetic datasets and consider more fonts, backgrounds, and varied noise conditions to create more realistic simulations.
@@
The analysis compared the performance of various OCR engines on Tamil and Sinhala printed scripts. Although fine-tuning the systems has the potential to improve performance, it was not carried out as this study intended only to analyse existing systems. It is also important to note a limitation regarding the datasets used; both are synthetically created, featuring black text on a white background. This design results in clean and clear images, which do not accurately reflect real-world conditions when capturing printed text. The decision to use synthetic data was made because similar realistic datasets were not available for the two low-resourced languages, and it was necessary to ensure a fair comparison.
While a considerable amount of work exists for Indic languages such as Hindi, according to the observations of de Silva (2025), the Devanagari script of Hindi has a distance of 5 from Sinhala and a distance of 7 from Tamil, as opposed to a distance of 4 between Sinhala and Tamil. Further, according to Rao (2021), scripts such as Sinhala and Tamil are considered rounded scripts, while Hindi is not. For these reasons, our work could not rely much on the progress made for Hindi OCR.
With real-world data, OCR technology faces significant challenges due to input image quality, particularly with historical documents and low-resource language data. Factors such as poor print quality, low resolution, shading, blurriness, and distortion can severely impact accuracy Hegghammer (2022). Images that exhibit distortion, textured backgrounds, cluttered environments, disconnected line segments, isolated dots, breaks in lines, rotation, motion blur, and out-of-focus blur complicate character segmentation and recognition, often leading to higher error rates Karunarathne et al. (2017). Additionally, low image resolutions can slow down OCR speed due to greater uncertainty in character representation Anuradha et al. (2021). Consequently, when evaluated using camera-captured images, the accuracy levels may notably differ from the results presented in this study.